OpenStudy (anonymous):

4(2m-n) -3(m-n) m= -15 , n= -18 I worked this and got 144 but my choices are 3 -3 -12 16 im confused?

OpenStudy (anonymous):

@Jhannybean

OpenStudy (jhannybean):

first expand the function

OpenStudy (jhannybean):

4(2m) -4n -3m + 3n = 8m -4n -3m +3n = (8 - 3)m + (-4 + 3)n

OpenStudy (jhannybean):

Simplify that.

OpenStudy (jhannybean):

then just substitute your values of m and n, you should get the right answer

OpenStudy (jhannybean):

Ive got to head off, goo luck!

OpenStudy (anonymous):

Are you still confused?

OpenStudy (anonymous):

uh yeah just a tad..